OpenStudy (anonymous):

whats the vertex of q (x)=x^2+6x+5

jhonyy9 (jhonyy9):

(x+1)(x+5)

OpenStudy (anonymous):

to find the vertex use \[x=-\frac{b}{2a}\] for the first coordinate

OpenStudy (anonymous):

in this case you have \[q(x)=x^2+6x+5\] so \[-\frac{b}{2a}=-\frac{6}{2}=-3\] for the first coordinate of the vertex. replace x by -3 to find the second coordinate

OpenStudy (anonymous):

q (x)=x^2+6x+5 =x^2+2*x*3 + 9 -4 =(x-3)^2-4 =a(x-h)^2-k vertex(h,k)=(-3,4)